# Patterns of imbalance states between sub-brain regimes during development in the resting state

### v1

@@ Abstract
The functional brain network is the result of neurons and different areas working together in a coordinated manner. The findings indicate that the spatial distribution of anticorrelations in the brain network is not inherent. In this study, using Heider balance theory, we examine the network of imbalanced triangles with negative links in the brain, which undergoes dynamic changes across development. By analyzing the age-related transformations in the connectivity of imbalanced triangles at the level of brain subnetworks. We show that anticorrelations evolve from childhood to adulthood, reflecting a developmental trajectory from a locally modular organization in childhood, through a flexible and reconfigurable architecture during adolescence, to a highly segregated and functionally specialized network system in adulthood. This mature configuration supports a balance between internally directed cognition and external task demands, enabling efficient and adaptive cognitive control. Specifically, in childhood, prominent anticorrelations were observed between the default mode network (DMN) and the dorsal attention network (DAN), as well as between the default mode network (DMN) and the ventral attention network (VAN). During adolescence, these anticorrelations persisted but became weaker. By adulthood, dominant anticorrelations were observed between the DMN and both the DAN and the frontoparietal network (FPN).
Keywords fMRI  $\cdot$ Heider Balance  $\cdot$ sub-brain regimes  $\cdot$ resting state
@@ Introduction
The application of graph theory and network science to the study of functional brain connectivity was pioneered by researchers such as Olaf Sporns sporns2004motifs , who introduced graph-theoretical methods to map and analyze the brain’s functional connectivity. Other seminal contributions were made by Edward Bullmore hagmann2008mapping , who highlighted the small-world properties of brain networks, and Michael G. Watts and Steven H. Strogatz watts1998collective , whose small-world network model provided the foundation for understanding the efficiency and modularity of brain connectivity. Numerous studies have investigated the brain using network science approaches. This framework has been widely applied in both healthy and clinical populations, offering insights into typical development, aging, and neurological disorders.dickerson2009large ; cohen2020brain ; haqiqifar2025patterns Recent studies have applied structural balance theory (SBT) to brain networks, revealing that complex interactions beyond pairwise connections significantly shape their dynamics.
@@
[…]
    • A batch vector (215×1215$\times$ 1215 × 1) was created to represent site codes (from 1 to 5).
    • A model matrix (215×2215$\times$ 2215 × 2) containing age and group labels (all healthy controls) was used to preserve biological variation during harmonization.
@@
The concept of balance theory was first introduced by Heider in 1946 in the field of social psychology. The theory considers triadic relationships; in other words, the theory states that when two individuals are in a friendship, the introduction of a third person who is in a relationship with both of them has an impact on the relationship between the two. Heider assigned positive (negative) sign value to pairwise interactions in the context of friendship (enmity) relationships. A balanced (imbalanced) relationship is defined as one in which the product of the signs is positive (negative). heider1946attitudes ; heider2013psychology In this context, four types of triads are defined: balanced ($Delta_{+++}$), balanced ($Delta_{+--}$), imbalanced ($Delta_{++-}$), and imbalanced ($Delta_{---}$) Fig.1.
Figure 1: Typical balanced and imbalanced triangles
@@
A fundamental tenet of the balance theory is that certain structures are in a balanced state, while others are not heider2013psychology ; wasserman1994social . It is important to note that an imbalanced triad represents a frustrated configuration, as it creates tension within the network and drives its links to adjust in a way that resolves the imbalance and restores a balanced state saberi2021topological . In the real world, people are involved in many tripartite relationships at the same time, so balance theory should apply to the set. This work was first done by Cartwright and Harari. They developed the notion into the field of graph theory, which was subsequently named the Structural Balance Theory cartwright1956structural . Heider’s theory only refers to the psychological structure of people, but by extending this theory, it made it possible to use this theory for the study of various social, biological, economic, and… networks. Inspired by the graph-theoretical framework, Structural Balance Theory has been developed as an approach to analyzing the organizational properties of complex networks antal2006social . While graph theory was initially formulated to model dyadic relationships between information units, SBT extends beyond pairwise interactions to examine triadic relationships, offering deeper insights into network structures.moradimanesh2021altered When applying graph-theoretical analysis to brain networks, a graph G(V, E) is defined, where V represents a set of nodes corresponding to specific brain regions. The strength of association between each pair of regions of interest is then quantified as a set of weighted links (E), providing insight into the functional connectivity of the brain. In Structural Balance Theory, the energy of a network ($E(N)$) measures its balance. The balance energy is calculated as shown in Equation 1. A positive product of edge signs within a triad indicates a balanced interaction, while a negative product signals imbalance. The energy is computed by summing the negative of these products across all triads in the network and normalizing by the total number of possible triads. Specifically, for a network with N nodes, the total number of possible triads is given by the combination formula $\binom{N}{3}$ This normalization ensures that $E(N)$ is bounded within the range of 1-1- 1 to +1+1+ 1, where $E(N)=-1$ corresponds to a fully balanced network and $E(N)=+1$ indicates a fully imbalanced one. The negative sign in the summation reflects the analogy to physical energy, where higher energy values are associated with greater frustration, and lower energy values correspond to more stable, balanced configurations.
@@
As highlighted in the introduction, the brain operates as a complex network, where not only individual regions but also the interconnections between them are essential for comprehending its functional and structural dynamics. The Probability Density Function of correlations across 400 brain regions is illustrated by age groups in Fig.2a. The KL divergence values—Children to Adolescents: 0.015, Children to adults: 0.013, and Adolescents to adults: 0.018—indicate minimal differences across age groups, suggesting that the changes in correlation distributions are not substantial or noticeable. The heatmaps illustrate the correlation between brain regions across different age groups—children, adolescents, and adults—highlighting key patterns in functional connectivity. (see Fig.2b,Fig.2c,Fig.2d) Symmetry in these correlations, particularly between regions in the left and the right hemispheres, indicates balanced inter-hemispheric communication.
@@
(c) Adolescents
(d) Adults
Figure 2: a) Probability Density Function of correlation across brain regions by age groups and correlation heatmap for different age groups: b) Children, c) Adolescents, d) Adults.
In the previous section, we examined the correlation probability distribution function (PDF) across different brain regions, as well as the functional connectivity map of 400 brain regions. To investigate whether anti-synchronies are a global or local phenomenon, within the framework of Structural Balance Theory, we focus on frustrated triads in which all three links are negative. This is because in imbalanced triangles—where only one link is negative—it is not possible to observe the spread of anti-synchrony across all three regions simultaneously. In such cases, the single negative link induces anti-synchrony only between two regions, while the third region remains synchronized with the others. In contrast, in imbalanced triangles (with three negative links), we can simultaneously observe the side effects of anti-synchrony: when anti-synchrony arises between two regions, it also appears in their interactions with the third region. Therefore, by examining the connectivity network of these fully negative triads, we can trace the spread of anti-synchrony and observe how it evolves with aging. We first analyzed imbalanced triads with negative links distribution function across different developmental stages. The distribution is plotted on a semi-logarithmic scale to capture variations across orders of magnitude more effectively Fig.3. The KL divergence values for these comparisons are as follows: children vs. adolescents: 0.041, children vs. adults: 0.001, and adolescents vs. adults: 0.020.
Figure 3: Semi-log distribution function of imbalanced triangles with negative links by age group
@@
The results indicate that while the overall distribution function remains consistent across age groups, there are variations in the intercepts. As we can see, the number of imbalanced triangles increases as the balance energy approaches zero for each group. However, imbalanced triangles, which have lower balance energy, are noticeably fewer compared to those with small amounts of imbalance energy. We focus on imbalanced triangles whose balance energy is lower than the average observed across groups (i.e., less than 0.002-0.002- 0.002). Hereafter, we refer to the Below-Average Imbalanced Triangle as $\Delta_{---}$.
$\text{Mean of }\log\text{-scaled PDF ( Imbalanced Triangles)}_{\text{children}}$ =0.00151=-0.00151= - 0.00151
$\text{Mean of }\log\text{-scaled PDF ( Imbalanced Triangles)}_{\text{adolescents}}$ =0.00229=-0.00229= - 0.00229
$\text{Mean of }\log\text{-scaled PDF ( Imbalanced Triangles)}_{\text{adults}}$ =0.00194=-0.00194= - 0.00194
We further compute the exact number of imbalanced triangles and the exact number of $\Delta_{---}$ for each group. Additionally, we identify the subset of these triangles that establish connections with one another  [^1] By $\Delta_{---}$, which establishes connections with one another, we mean sharing a common edge between two $\Delta_{---}$., as summarized in Tab.1
Age Group Imbalanced Triangles $\Delta_{---}$ Identified $\Delta_{---}$ with Connections
Children 464830 86 81
Adolescents 461372 60 57
adults 704571 379 358
Table 1: Number of all the imbalanced triangles and the number of $\Delta_{---}$, and the number of $\Delta_{---}$ including those that establish direct connections, across developmental groups.
To answer whether $\Delta_{---}$ is a local or global phenomenon, we first examine their dendrograms. The modular structures of their network are visible in the dendrograms.
@@
The dendrogram analysis reveals distinct developmental patterns in the organization of $\Delta_{---}$ across age groups.(see Fig.4a,Fig.4b,Fig.4c) In adults, the network is significantly larger, with a high density of $\Delta_{---}$ distributed throughout, suggesting a more integrated but widespread presence of imbalances. The hierarchical clustering appears more continuous, with fewer clearly defined modules, indicating a more interconnected structure. In contrast, adolescence exhibits the fewest $\Delta_{---}$, with the largest group of imbalanced triangles more distinctly separated from the others, suggesting a transitional phase in which the brain is reorganizing its network. This may reflect developmental processes such as synaptic pruning or shifts in functional connectivity, reducing local imbalances before integrating them into a more efficient and interconnected adult structure. Childhood presents a higher number of triangles than adolescence, with strong modularity still visible, implying that imbalances are more confined within specific subsystems. The differences in sparsity and modular structure across age groups suggest that while $\Delta_{---}$ are initially localized in childhood, they become more distributed and integrated into the broader network by adulthood, reflecting a shift in the brain’s functional organization over development.
@@
To analyze the connection network between $\Delta_{---}$, we consider only $\Delta_{---}$ that are connected. In this network, each node represents a $\Delta_{---}$, and a link between two nodes indicates a common edge between the corresponding $\Delta_{---}$. We assign each $\Delta_{---}$ a numerical identifier that corresponds to its node number in the network.
@@
To visualize the functional networks Fig.4d,Fig.4e,Fig.4f, we employed Gephi applicationbastian2009gephi ; gephi , an open-source platform for network analysis and graph-based data visualization. The layout was optimized using the Force Atlas-3D algorithm to enhance interpretability.
@@
Anti-correlation, or anti-synchrony, plays a crucial role in the organization and functionality of brain networks. In a foundational study by Fox et al. (2005)fox2005human , it was demonstrated that the brain is intrinsically organized into dynamic, anticorrelated functional networks, particularly between the default mode network (DMN) and task-positive networks. These anticorrelated patterns suggest that anti-synchrony is fundamental to the brain’s ability to switch between different cognitive states, thereby supporting efficient information processing and behavioral flexibility. In a more recent study, the importance of negative links was further emphasized in the context of emotional processing. The authors showed that negative connections contribute significantly to the structural balance and stability of brain networks during emotional experiences, highlighting their role in maintaining functional balance soleymani2023impact . Together, these findings underscore that negative or anticorrelated interactions are not mere noise but essential components of healthy brain dynamics. To investigate whether brain function driven by anticorrelations is a local or global phenomenon, we analyzed the network of ($\Delta_{---}$) based solely on their connectivity. The structure of these networks across different age groups reveals that the brain exhibits specific organizational patterns in how $\Delta_{---}$ interacts. An important question that arises from this network structure is whether the brain needs to coordinate and confirm all regional relationships to establish a connection, or whether it allows regions to act independently within local contexts?
@@
In childhood and adolescence, the $\Delta_{---}$ connection networks appear more discrete, indicating that brain functions operate more independently of one another during these developmental stages. The low KL divergence values between the distribution functions of $\Delta_{---}$ further support this interpretation, suggesting that there is no significant tendency for $\Delta_{---}$ to form connections with one another in these age groups. This lack of integration implies that decisions or functional changes within the brain are made locally, without the need for coordination or "permission" from the broader network.
@@
In adulthood, however, the pattern is markedly different. The number of giant components in the $\Delta_{---}$ connection networks is noticeably higher compared to those in childhood and adolescence, indicating that a greater number of $\Delta_{---}$ are interconnected. This suggests that, unlike in earlier stages, the brain’s functional organization in adulthood increasingly relies on the integration of imbalanced structures. During this developmental period, the function of the brain appears to emerge from a larger, more cohesive network of interacting $\Delta_{---}$, reflecting a shift toward global coordination and functional integration. The emergence of anti-synchrony between two brain regions does not affect only those regions—it can spread across the brain network, prompting a global response from the brain as a complex system. This propagation of anti-correlation can influence various parts of the system. Consider a societal analogy: an increase in the dollar exchange rate leads to a rise in pesticide prices. Eventually, the price of milk also increases—not because the dollar is directly linked to milk, but because the rising cost of pesticides affects livestock management. Farmers cannot afford sufficient pesticides, their feed production declines, and as a result, milk production drops, leading to higher milk prices. Similarly, in adulthood, the emergence of anti-synchrony between two brain regions does not remain localized. Other brain regions are also affected as part of the broader network response to this local disruption.
@@
In adulthood, brain function driven by the interactions between $\Delta_{---}$ triangles emerges from a larger and more integrated network. Cognitive processes during this stage require greater coordination, and decisions tend to be made at a global level. This reflects a form of harmony and synchronization between the $\Delta_{---}$ and overall brain activity. At times, the adult brain network may even tolerate a small frustration if it benefits the system as a whole. In other words, each decision is evaluated collectively, and a final outcome is reached only when it is confirmed across the broader network. In adulthood, the brain appears to adopt a specific communication pattern for managing anticorrelations. The interactions between $\Delta_{---}$ triangles serve as a framework for evaluating and guiding decision-making processes.
@@
Figure 4: Dendogram and Networks of ($\Delta_{---}$) across age groups: a,d) Children, b,e) Adolescents, c,f) Adults. It shows modularity in the triangle networks.
To identify which $\Delta_{---}$ are connected and the specific energy levels at which these connections occur, we analyze the energy-energy (e-e) graph. The energy-energy (e-e) graphs highlight key developmental differences in the distribution and integration of $\Delta_{---}$ across age groups(see Fig.5a,Fig.5b,Fig.5c). In childhood, the $\Delta_{---}$ triangles that share common edges tend to have energies within a narrow interval, specifically between 0.003-0.003- 0.003 and 0.002-0.002- 0.002, indicating that only higher-energy structures are functionally interconnected at this stage. A similar pattern is observed during adolescence, despite a lower overall number of $\Delta_{---}$. In contrast, adulthood exhibits a markedly different organization: there is a strong synchronization between $\Delta_{---}$ triangles across a broader energy spectrum. In this stage, both high- and low-energy $\Delta_{---}$ are functionally linked within the network, and their overall number is significantly increased. This suggests a more integrated and cohesive system in adulthood, where even imbalanced structures are incorporated into the broader network architecture.
@@
Figure 5: energy-energy of $\Delta_{---}$ plots for different age groups: a) Children, b) Adolescents, c) Adults.
@@
We now aim to examine the $\Delta_{---}$ participating in Networks Fig.4d,Fig.4e,Fig.4f from a different perspective, taking a step back to observe the broader context. Each node in Networks Fig.4d,Fig.4e,Fig.4f represents a $\Delta_{---}$, with each of its vertices corresponding to one of the 400 distinct brain regions. We are particularly interested in identifying which brain subnetworks these vertices belong to—specifically, the subnetworks that are involved in Networks Fig.4d,Fig.4e,Fig.4f. To answer this question, we analyze all the nodes in Networks Fig.4d,Fig.4e, Fig.4f. Since each node is a triangle, each of its edges connects two brain subnetworks through an anti-synchronous (anticorrelated) relationship. To visualize the anti-synchronous interactions between the subnetworks derived from Networks Fig.4d,Fig.4e,Fig.4f, we define a new network Fig.6. In this derived network, the nodes represent the brain subnetworks, and the links between them indicate the number of edges that connect the respective subnetworks negatively. Therefore, the higher the weight of a link, the more negative connections exist between the two subnetworks. These networks were constructed and visualized using Gephi application bastian2009gephi ; gephi . Tab.2,Tab.3, and Tab.4 summarize the connection weights between subnetworks, as well as the proportion of each weight relative to the total weight, for the children, adolescent, and adult groups, respectively. It is important to emphasize that this derived network is extracted specifically from Networks Fig.4d,Fig.4e,Fig.4f. This means that it does not reflect all anticorrelations present in the brain but only those that form the edges of $\Delta_{---}$ involved in the structure of Networks Fig.4d,Fig.4e,Fig.4f. The size of each node reflects its degree—that is, the number of connections it has—so larger nodes are more extensively involved in anticorrelated interactions. Furthermore, the color of each link transitions from red to blue as its weight decreases, with red indicating a stronger (more frequent) anticorrelation between subnetworks.

### v2

# Dynamic pattern of imbalance states between resting-state functional brain networks -during development

@@ Abstract
The functional brain network emerges from the complex, coordinated activity of distinct yet connected regions, which underlie the diverse repertoire of human cognitive functions. Structural Balance Theory (SBT) has been successfully applied to model such nontrivial connections through the analysis of balance and unbalance triadic configurations. In this study, using SBT, we examine the network of imbalanced triads in the resting-state brain subnetworks, which undergo dynamic changes during development. We demonstrate that anticorrelation patterns evolve across the lifespan, reflecting a developmental trajectory from a locally modular organization in childhood to a flexible and reconfigurable architecture during adolescence and finally to a highly segregated and functionally specialized network system in adulthood. This developmental trajectory indicates that the spread of anticorrelations is not an inherent feature of brain organization. This mature organization facilitates a balance between self-referential, internally generated cognitive processes and externally oriented, goal-directed cognition, enabling efficient and adaptive cognitive control. This balance is underpinned by prominent anticorrelations between the Default Mode Network (DMN) and the Frontoparietal Network (FPN) in adulthood. This is while during adolescence these anticorrelations are substantially weaker, suggesting that the maturation of these network connections from adolescence to adulthood establishes a functional architecture that supports the segregation of internal and external cognitive processes. These findings elucidate how the dynamic evolution of anticorrelation patterns in brain networks supports cognitive development across the lifespan, offering new insights into the neural basis of adaptive cognitive control.
Keywords Network Neuroscience $\cdot$ Structural Balance Theory (SBT)  $\cdot$ resting-state functional MRI (rs-fMRI)
@@ Introduction
The application of graph theory and network science to the study of functional brain connectivity was pioneered by researchers such as Olaf Sporns sporns2004motifs , who introduced graph-theoretical methods to map and analyze the brain’s functional connectivity. Other seminal contributions were made by Edward Bullmore hagmann2008mapping , who highlighted the small-world properties of brain networks, and Michael G. Watts and Steven H. Strogatz watts1998collective , whose small-world network model provided the foundation for understanding the efficiency and modularity of brain connectivity. Numerous studies have investigated the brain using network science approaches. This framework has been widely applied in both healthy and clinical populations, offering insights into typical development, aging, and neurological disorders dickerson2009large ; cohen2020brain ; haqiqifar2025patterns . Recent studies have applied structural balance theory (SBT) to brain networks, revealing that complex interactions beyond pairwise connections significantly shape their dynamics.
@@
    • For each participant, the lower triangle of the $400\times 400$ functional connectivity matrix was vectorized to form a $79{,}800\times 1$ feature vector.
    • These vectors were concatenated column-wise to produce a $79{,}800\times 215$ data matrix, where each column corresponds to one participant.
    • A batch vector ($215\times 1$) was created to represent site codes (from 1 to 5).
    • A model matrix ($215\times 2$) containing age and group labels (all healthy controls) was used to preserve biological variation during harmonization.
@@
The concept of balance theory was first introduced by Heider in 1946 in the field of social psychology. The theory considers triadic relationships; in other words, the theory states that when two individuals are in a friendship, the introduction of a third person who is in a relationship with both of them has an impact on the relationship between the two. Heider assigned positive (negative) sign value to pairwise interactions in the context of friendship (enmity) relationships. A balanced (imbalanced) relationship is defined as one in which the product of the signs is positive (negative). heider1946attitudes ; heider2013psychology In this context, four types of triads are defined: balanced ($\Delta_{+++}$), balanced ($\Delta_{+--}$), imbalanced ($\Delta_{++-}$), and imbalanced ($\Delta_{---}$) Fig.1.
Figure 1: Illustration of the four possible triad types defined by Structural Balance Theory (SBT): balanced ($\Delta_{+++}$), balanced ($\Delta_{+--}$), imbalanced $\Delta_{++-}$, and fully imbalanced $\Delta_{---}$
@@
A fundamental tenet of the balance theory is that certain structures are in a balanced state, while others are not heider2013psychology ; wasserman1994social . It is important to note that an imbalanced triad represents a frustrated configuration, as it creates tension within the network and drives its links to adjust in a way that resolves the imbalance and restores a balanced state saberi2021topological . In the real world, people are involved in many tripartite relationships at the same time, so balance theory should apply to the set. This work was first done by Cartwright and Harari. They developed the notion into the field of graph theory, which was subsequently named the Structural Balance Theory (SBT) cartwright1956structural . Heider’s theory only refers to the psychological structure of people, but by extending this theory, it made it possible to use this theory for the study of various social, biological, economic, and… networks. Inspired by the graph-theoretical framework, SBT has been developed as an approach to analyzing the organizational properties of complex networks antal2006social . While graph theory was initially formulated to model dyadic relationships between information units, SBT extends beyond pairwise interactions to examine triadic relationships, offering deeper insights into network structures.moradimanesh2021altered When applying graph-theoretical analysis to brain networks, a graph G(V, E) is defined, where V represents a set of nodes corresponding to specific brain regions. The strength of association between each pair of regions of interest is then quantified as a set of weighted links (E), providing insight into the functional connectivity of the brain. In SBT, the energy of a network ($E(N)$) measures its balance. The balance energy is calculated as shown in Equation 1. A positive product of edge signs within a triad indicates a balanced interaction, while a negative product signals imbalance. The energy is computed by summing the negative of these products across all triads in the network and normalizing by the total number of possible triads. Specifically, for a network with N nodes, the total number of possible triads is given by the combination formula $\binom{N}{3}$ . This normalization ensures that $E(N)$ is bounded within the range of 1-1 to +1+1, where E(N)=1E(N)=-1 corresponds to a fully balanced network and $E(N)=+1$ indicates a fully imbalanced one. The negative sign in the summation reflects the analogy to physical energy, where higher energy values are associated with greater frustration, and lower energy values correspond to more stable, balanced configurations.
@@
Identification of Connected Imbalances Triads
To investigate whether anti-synchronies are a global or local phenomenon, within the framework of SBT, we focus on frustrated triads in which all three links are negative. In other words, we focus on triple imbalances that reflect antisynchronous relationships between three regions of the brain that were obtained from the 400×400 brain network constructed in our work. This is because in imbalanced triads, where only one link is negative, it is not possible to observe the spread of antisynchrony across all three regions simultaneously. In such cases, the single negative link induces anti-synchrony only between two regions, while the third region remains synchronized with the others. In contrast, in imbalanced triads (with three negative links), we can simultaneously observe the side effects of anti-synchrony: when anti-synchrony arises between two regions, it also appears in their interactions with the third region. Therefore, by examining the connectivity network of these fully negative triads, we can trace the spread of anti-synchrony and observe how it evolves with aging. We considered those triads whose balance energy exceeded the average balance energy of all $\Delta_{---}$ as $\Delta{{}^{+}}_{---}$. In addition, we restricted our analysis to cases in which such triads were connected, meaning that two $\Delta_{---}$ shared a common edge. We denote these connected triads as $\Delta{{}^{*}}_{---}$.
@@
Whole-Brain Functional Correlation and Developmental Patterns
As highlighted in the introduction, the brain operates as a complex network, where not only individual regions but also the interconnections between them are essential for comprehending its functional and structural dynamics. The Probability Density Function of correlations across 400 brain regions is illustrated by age groups in Fig.2a. The KL divergence values—Children to Adolescents: 0.015, Children to adults: 0.013, and Adolescents to adults: 0.018—indicate minimal differences across age groups, suggesting that the changes in correlation distributions are not substantial or noticeable. The heatmaps illustrate the correlation between brain regions. Due to the minimal differences between the KL divergence values across age groups and the absence of essential changes in the correlation distributions, we present only the heatmap for adults. (see Fig.2b) Symmetry in these correlations, particularly between regions in the left and the right hemispheres, indicates balanced inter-hemispheric communication.
@@
(b) Adults
Figure 2: (a) Probability Density Function of correlations among 400 brain regions across age groups. (b) Heatmaps of functional correlations for adults, showing subtle developmental refinements while maintaining overall stability in correlation structure.
The Distribution and Abundance of $\Delta_{---}$
In the previous section, we examined the correlation Probability Distribution Function (PDF) across different brain regions, as well as the functional connectivity map of 400 brain regions. In this section we analyzed $\Delta_{---}$ the distribution function across different developmental stages. The distribution is plotted on a semi-logarithmic scale to capture variations across orders of magnitude more effectively Fig.3b. The KL divergence values for these comparisons are as follows: children vs. adolescents: 0.041, children vs. adults: 0.001, and adolescents vs. adults: 0.020.
(a) Distribution Function of $\Delta_{---}$ by age groups
@@
Figure 3: a) Distribution function of $\Delta_{---}$ for children, adolescents, and adults. b) Exact counts of $\Delta_{---}$, $\Delta{{}^{+}}_{---}$, and $\Delta{{}^{*}}_{---}$ for each age group, highlighting developmental differences in balance energy profiles.
@@
The results indicate that while the overall distribution function remains consistent across age groups, there are variations in the intercepts. As illustrated in Fig. 3a, the number of $\Delta_{---}$ increases as the balance energy approaches zero across all age groups. However, triads with lower balance energy are markedly fewer compared to those with small imbalance energy, which we previously denoted as $\Delta^{+}_{---}$. In our analysis, we focus on ($\Delta^{*}_{---}$) as we said before. Specifically, the mean balance energy of the Distribution Function of $\Delta_{---}$ was 0.00151-0.00151 for children, 0.00229-0.00229 for adolescents, and 0.00194-0.00194 for adults; therefore, we selected connected $\Delta_{---}$ with balanced energy less than 0.002-0.002 as $\Delta^{*}_{---}$. Figure 3b presents the exact numbers of $\Delta_{---}$, $\Delta^{+}{---}$, and $\Delta^{+}{---}$ for each group.
Modularity and Network Organization of $\Delta^{*}_{---}$
@@
To answer whether $\Delta{{}^{+}}_{---}$ is a local or global phenomenon, we first examine its dendrograms. The modular structures of their network are visible in the dendrograms (see Fig.4a,4b,4c). The dendrogram analysis reveals distinct developmental patterns in the organization of $\Delta{{}^{*}}_{---}$ across age groups. In adults, the network is significantly larger, with a high density of $\Delta{{}^{+}}_{---}$ distributed throughout, suggesting a more integrated but widespread presence of imbalances. The hierarchical clustering appears more continuous, with fewer clearly defined modules, indicating a more interconnected structure. In contrast, adolescence exhibits the fewest $\Delta{{}^{+}}_{---}$, with the largest group of imbalanced triads more distinctly separated from the others, suggesting a transitional phase in which the brain is reorganizing its network. This may reflect developmental processes such as synaptic pruning or shifts in functional connectivity, reducing local imbalances before integrating them into a more efficient and interconnected adult structure. Childhood presents a higher number of triads than adolescence, with strong modularity still visible, implying that imbalances are more confined within specific subsystems. The differences in sparsity and modular structure across age groups suggest that while $\Delta{{}^{+}}_{---}$ are initially localized in childhood, they become more distributed and integrated into the broader network by adulthood, reflecting a shift in the brain’s functional organization over development.
@@
To analyze the connection network between $\Delta{{}^{+}}_{---}$,as we mentioned we consider only $\Delta{{}^{+}}_{---}$ that are connected. In this network, each node represents a $\Delta{{}^{*}}_{---}$, and a link between two nodes indicates a common edge between the corresponding $\Delta{{}^{*}}_{---}$. We assign each $\Delta{{}^{*}}_{---}$ a numerical identifier that corresponds to its node number in the network (see Fig.4a,4e,4f).
@@
To visualize the functional networks we employed Gephi applicationbastian2009gephi ; gephi , an open-source platform for network analysis and graph-based data visualization. The layout was optimized using the Force Atlas-3D algorithm to enhance interpretability.
@@
Anti-correlation, or anti-synchrony, plays a crucial role in the organization and functionality of brain networks. In a foundational study by Fox et al. (2005)fox2005human , it was demonstrated that the brain is intrinsically organized into dynamic, anticorrelated functional networks, particularly between the Default Mode Network (DMN) and task-positive networks. These anticorrelated patterns suggest that anti-synchrony is fundamental to the brain’s ability to switch between different cognitive states, thereby supporting efficient information processing and behavioral flexibility. In a more recent study, the importance of negative links was further emphasized in the context of emotional processing. The authors showed that negative connections contribute significantly to the structural balance and stability of brain networks during emotional experiences, highlighting their role in maintaining functional balance soleymani2023impact . Together, these findings underscore that negative or anticorrelated interactions are not mere noise but essential components of healthy brain dynamics. To investigate whether brain function driven by anticorrelations is a local or global phenomenon, we analyzed the network of $\Delta{{}^{*}}_{---}$. The structure of these networks across different age groups reveals that the brain exhibits specific organizational patterns in how $\Delta{{}^{*}}_{---}$ interacts. An important question that arises from this network structure is whether the brain needs to coordinate and confirm all regional relationships to establish a connection, or whether it allows regions to act independently within local contexts.
@@
In childhood and adolescence, the $\Delta{{}^{*}}_{---}$ networks appear more discrete, indicating that brain functions operate more independently of one another during these developmental stages. The low KL divergence values between the distribution functions of $\Delta_{---}$ further support this interpretation, suggesting that there is no significant tendency for $\Delta_{---}$ to form connections with one another in these age groups. This lack of integration implies that decisions or functional changes within the brain are made locally, without the need for coordination or "permission" from the broader network.
@@
In adulthood, however, the pattern is markedly different. The number of giant components in the $\Delta{{}^{*}}_{---}$ networks is noticeably higher compared to those in childhood and adolescence, indicating that a greater number of $\Delta{{}^{*}}_{---}$ are interconnected. This suggests that, unlike in earlier stages, the brain’s functional organization in adulthood increasingly relies on the integration of imbalanced structures. During this developmental period, the function of the brain appears to emerge from a larger, more cohesive network of interacting $\Delta{{}^{*}}_{---}$, reflecting a shift toward global coordination and functional integration. The emergence of anti-synchrony between two brain regions does not affect only those regions—it can spread across the brain network, prompting a global response from the brain as a complex system. This propagation of anti-correlation can influence various parts of the system. Consider a societal analogy: an increase in the dollar exchange rate leads to a rise in pesticide prices. Eventually, the price of milk also increases—not because the dollar is directly linked to milk, but because the rising cost of pesticides affects livestock management. Farmers cannot afford sufficient pesticides, their feed production declines, and as a result, milk production drops, leading to higher milk prices. Similarly, in adulthood, the emergence of anti-synchrony between two brain regions does not remain localized. Other brain regions are also affected as part of the broader network response to this local disruption.
@@
In adulthood, brain function driven by the interactions between $\Delta{{}^{*}}_{---}$ triads emerges from a larger and more integrated network. Cognitive processes during this stage require greater coordination, and decisions tend to be made at a global level. This reflects a form of harmony and synchronization between the $\Delta{{}^{*}}_{---}$ and overall brain activity. At times, the adult brain network may even tolerate a small frustration if it benefits the system as a whole. In other words, each decision is evaluated collectively, and a final outcome is reached only when it is confirmed across the broader network. In adulthood, the brain appears to adopt a specific communication pattern for managing anticorrelations. The interactions between $\Delta{{}^{*}}_{---}$ triads serve as a framework for evaluating and guiding decision-making processes.
Energy–Energy Connectivity of $\Delta{{}^{*}}_{---}$
To identify which $\Delta{{}^{*}}_{---}$ are connected and the specific energy levels at which these connections occur, we analyze the energy-energy (e-e) graph. The energy-energy (e-e) graphs highlight key developmental differences in the distribution and integration of $\Delta{{}^{*}}_{---}$ across age (see Fig.5). In childhood, the $\Delta{{}^{*}}_{---}$ that share common edges tend to have energies within a narrow interval, specifically between 0.003-0.003 and 0.002-0.002, indicating that only higher-energy structures are functionally interconnected at this stage. A similar pattern is observed during adolescence, despite a lower overall number of $\Delta{{}^{*}}_{---}$. In contrast, adulthood exhibits a markedly different organization: there is a strong synchronization between $\Delta{{}^{*}}_{---}$ across a broader energy spectrum. In this stage, both high- and low-energy $\Delta{{}^{*}}_{---}$ are functionally linked within the network, and their overall number is significantly increased. This suggests a more integrated and cohesive system in adulthood, where even imbalanced structures are incorporated into the broader network architecture.
@@
Figure 4: (a–c) Dendrograms of $\Delta{{}^{*}}_{---}$ in children, adolescents, and adults, respectively. (d–f) Corresponding network visualizations reveal modularity and integration patterns, showing that $\Delta{{}^{*}}_{---}$ shifts from localized clusters in childhood to more interconnected structures in adulthood.
@@
Figure 5: (a–c) energy–energy plots for children, adolescents, and adults. Results show that while only higher-energy imbalanced triads connect in childhood and adolescence, adulthood exhibits broad-spectrum connectivity, suggesting increased global integration of anticorrelations.
Subnetwork-Level Anticorrelations Emerging from $\Delta{{}^{*}}_{---}$
@@
We now aim to examine the $\Delta{{}^{*}}_{---}$ participating in Networks Fig.4d,Fig.4e,Fig.4f from a different perspective, taking a step back to observe the broader context. Each node in these Networks represents a $\Delta{{}^{*}}_{---}$, with each of its vertices corresponding to one of the 400 distinct brain regions. We are particularly interested in identifying which brain subnetworks these vertices belong to—specifically, the subnetworks that are involved in the Networks. To answer this question, we analyze all the nodes in the Networks. Since each node is a triad, each of its edges connects two brain subnetworks through an anti-synchronous (anticorrelated) relationship. To visualize the anti-synchronous interactions between the subnetworks, we define a new network, Fig.6. In this derived network, the nodes represent the brain subnetworks, and the links between them indicate the number of edges that connect the respective subnetworks negatively. Therefore, the higher the weight of a link, the more negative connections exist between the two subnetworks. Tab.1, Tab.2, and Tab.3 present the connections with the highest weights between subnetworks, highlighting the links with considerable contributions to the overall network, for the children, adolescent, and adult groups, respectively. It is important to emphasize that this derived network is extracted specifically from Networks Fig.4d,Fig.4e,Fig.4f. This means that it does not reflect all anticorrelations present in the brain but only those that form the edges of $\Delta{{}^{*}}_{---}$.